# Cheick Tidiane Ba, Richard G. Clegg

### v1

@@
Abstract
In the dynamic landscape of the Web, we are witnessing the emergence of the Web3 paradigm, which dictates that platforms should rely on blockchain technology and cryptocurrencies to sustain themselves and their profitability. Web3 has garnered significant attention as the currencies gained considerable monetary value. Cryptocurrencies are characterised by high market volatility and susceptibility to substantial crashes, issues that require temporal analysis methodologies able to tackle the high temporal resolution, heterogeneity and scale of blockchain data. While existing research attempts to analyse crash events, fundamental questions persist regarding the optimal time scale for analysis, differentiation between long-term and short-term trends, and the identification and characterisation of shock events within these decentralised systems.
@@ Abstract
This paper pioneers temporal, cross-chain graph analysis to explore a cryptocurrency collapse. It emphasises the importance of temporal analysis for studies on web-derived data. In addition, the methodology shows how graph-based analysis can enhance traditional econometric results. Overall, this research carries implications beyond its field, for example for regulatory agencies aiming to safeguard users from shocks and monitor investment risks for citizens and clients.
Keywords transaction network  $\cdot$ temporal network  $\cdot$ cryptocurrency  $\cdot$ Web3  $\cdot$ multilayer graph
1 Introduction
@@
Analysing multi-layer, temporal networks at scale is an important problem in the study of complex networks. In particular, there are issues of understanding the best time-scale to analyse, how to determine long-term and short-term trends and how to recognise and characterise shock events in such a system. In this paper, we look at Web3 and specifically into six cryptocurrencies traded over the Ethereum blockchain. We use multi-layer, temporal network analysis techniques enabled by the Raphtory software developed at Queen Mary University of London and Pometry Ltd. The software enables simultaneously looking at how temporal networks evolve in time and at different time-scales while considering several network layers together.
@@
Digital currencies remain an important but potentially volatile market. In particular, there is the possibility that a particular currency will completely crash, wiping the value of that coin to a tiny fraction of its previous trade value: these events can destroy investments and savings, impacting the lives of investors. It is, therefore, important to analyse the events that lead up to a crash and to look at whether the effects of a crash can destabilise other currencies. “Stablecoins" are an attempt to make cryptocurrencies that can be used for trading rather than speculation by “pegging" the value of the coin to a fixed exchange rate (often one dollar to one stablecoin). The stablecoin TerraUSD (UST) [^1] Throughout this paper we will refer to currency by the shortened name it was commonly traded under. was pegged to trade for \$1 and was associated with a companion currency LUNA that was intended to stabilise it. Over a few days in May 2022 both UST and LUNA had collapsed and never regained their value. The collapse is believed to have been precipitated by deliberate and coordinated large-scale sales of UST in an attempt to crash UST, drive down the price of Bitcoin (BTC) and derive profit from this.
@@
In this paper, we use techniques from complex network analysis and, in particular, temporal multi-layer analysis to look at the events leading up to and immediately following the collapse of LUNA and UST. UST and LUNA were traded on a cryptocurrency platform known as Terra and this data is not available to us. However, the two currencies were also traded in a “wrapped" form on the Ethereum platform, USTC (US Terra classic) and WLUNC (wrapped Luna classic) respectively. From the available set of Ethereum trades, we look at USTC and WLUNC trades and compare them with the other most highly traded stablecoins on the platform. This gives us a pool of six currencies trading over the same platform that can be compared as they evolve in the weeks before and after the collapse, hence the analysis considers 26,209,860 transactions occurring in a symmetrical analysis period around the day of the crash.
@@
To our knowledge, this is the first paper to use temporal, cross-chain, graph analysis to look at the collapse of a cryptocurrency. The analysis in this paper can only be achieved with multi-layer temporal graph systems as it is necessary to look at correlations between different graph layers and to consider the evolution of the system at different time scales. We show that the six currencies are initially highly correlated in terms of the number of sales but, post-crash, separate into two groups where the “failed" currencies correlate with each other but not the other four. We show that two high-volume selling events on UST had a significant effect on the structure of sales volumes before the crash (one event exactly coinciding with the crash) effectively decoupling its sales volume from all other currencies. We also characterise a significant “rally" event that temporarily caused large changes to the UST and LUNA ecosystems shortly after the collapse. The collapse event has significant effects on graph metrics such as graph density and average degree but also effects on measures like clustering coefficient and weakly connected components.
2 Background
@@
Stablecoins are digital currencies that attempt to peg their value to some other reference asset, for example the US dollar or a fixed amount of gold [1]. This is designed to reduce the volatility of the currency as the high volatility of some currencies can put users off adopting them for use as currency (as opposed to speculation). Many different stablecoins exist traded over several different blockchain systems. Digital currencies can be traded between different blockchains by a process known as “wrapping". In this process, a currency token originating on chain A is held in place and a wrapped coin representing the same token is minted on chain B [2]. This allows, for example, BTC to be spent on the Ethereum chain as Wrapped Bitcoin (WBTC).
@@
Table 1: Currencies under study, the acronym is the label that the currency uses for trading.
In this paper, we wish to study the ecosystem around the currencies Terra and Luna which ran on their own blockchain known as Terra. TerraUSD(UST) was a stablecoin pegged at one US dollar. Luna (LUNA) was a cryptocurrency designed to ensure the stability of the Terra stablecoin. This was achieved by an algorithmic process known as “Anchor", that automatically would swap Terra with Luna currency to adjust their prices: in fact, the protocol kept the UST exchange rate stable by varying the pool of Terra available [3], generating or destroying tokens when needed. The Terra blockchain data is unavailable to us, however, we can study wrapped versions of UST and LUNA traded on the Ethereum blockchain. We provided a summary of the currencies under study in Table 1.
The interest in studying the Terra/Luna ecosystem comes from the sudden complete crash of UST in May 2022. A timeline extracted from [3] and [4] is as follows:
@@
  • 7th May 2022: Agreement is made to buy a large number of UST for BTC greatly reducing the liquidity in the pool. Simultaneously a large number of UST tokens are put up for sale (which caused the first UST de-pegging, below \$0.99 [3]).
@@
  • 25th May 2022: Proposal [5, 6, 7] for the launch of a new blockchain Terra 2.0 is official.
  • 27th May 2022 (T2): Terra 2.0 is live. Original LUNA tokens became Luna Classic (LUNC). The original UST stablecoin became TerraClassicUSD (USTC). Unlike the original Terra network, Terra 2.0 has no algorithmic stablecoin [8]. We denote this event as Terra 2.0 (T2).
The final consequence was not only the collapse of the Terra platform, UST and LUNA but also a decrease in the price of BTC. The motivation speculated for the attack is that the attackers were short-selling large amounts of BTC. Short selling is, in effect, a bet that a traded asset will decrease in price. The short seller contracts to borrow and sell the asset now and buy it at a later date at market price. If the price drops the seller, has profited by the amount of the price drop. In some markets, short selling is highly regulated as it has been implicated in market crashes [9]. In our case study, there is a belief that high levels of selling of Bitcoin observed in the days prior to the crisis were linked to short selling  [3].
@@
The Luna crash serves as an interesting case study due to its relevance within the stablecoins ecosystem, attracting attention from both industry and the research community. There are many works describing the ecosystem and the properties of the Terra system. For instance, Uhlig [10] proposes a theoretical framework for the analysis of hourly data on prices and market capitalisation data, delineating the gradual unfolding of the crash. Limited research leverages on-chain data. Cho [11] investigates the underlying causes of the de-pegging event focusing on the token economics of the Terra Blockchain. They utilise on-chain data focusing on transactions converting USTC to LUNC over the four days following the May 9 crash, identifying critical design flaws in the conversion process exploited by attackers. These works focus mainly on Terra. However, various studies explore the impact of the crash on other currencies. De Blasis et al. [12] examine the impact of the USTC collapse on other stablecoins by using proprietary minute-by-minute price transaction data of various cryptocurrencies, over a symmetrical 40-day period around the USTC crash of May 9th. Their work shows the presence of significant contagion from the USTC collapse, which could be explained by traders’ herding behaviour i.e. users reacting to the crash by moving to other currencies. Briola et al. [3] provide a background on the Terra project and describe the main events that led to the crash, before delving into an analysis focused on the hourly price data time series. They study the evolution of dependencies between 61 cryptocurrencies, by leveraging the price time series to define a network structure among currencies, a correlation network: their analysis seems to exclude the presence of herding behaviour. The importance of analyzing cross-chain behaviour is denoted by other works as well. Barthere et al. [13] analyses proprietary data from Terra and Ethereum to identify the main users involved in the crash. They assert that a small subset of users initiated the crash by leveraging system vulnerabilities, benefiting significantly by converting USTC to other coins before the crash. Liu et al. [14] employ a similar dataset and approach, leveraging proprietary transaction data from Terra and Ethereum. Their extensive analysis across multiple chains and assets reveals the complex nature of the run on Terra. At the same time, they believe the crash was exacerbated by concerns about system sustainability and not just a concentrated market manipulation. Although they have user transaction data, in their study, the authors do not focus on the transaction networks but focus more on describing the events and identifying the main actors.
@@
A key limitation in these works is while they provide important results on the crash, they neglect the analysis of the transactions among users from a network standpoint: indeed, the circulation of tokens within the ecosystem can be effectively represented as networks or graphs [15]. In this representation, nodes (or vertices) symbolise users, while links (or edges) denote the dynamic interactions or relationships among them. This modelling and analysis approach for currency systems through networks has significantly enhanced our understanding of financial transaction networks [16, 17]. Moreover, temporal networks [18] have proven to be a valuable framework for comprehending web-based currency systems such as Bitcoin [19, 20, 21, 22] and Ethereum [23, 24, 25, 26], as well as in the context of new-generation social media platforms [27, 28, 29]. However, none of these works analyses the stablecoin ecosystem from a network standpoint. Recent works analyzing the Chartalist dataset by Shamsi et al. [4] highlighted the importance of network structure in the stablecoin system. For instance, Huang et al. [30] focus on change detection on dynamic graphs, proposing a spectral method to create dynamic graph embeddings and an associated anomaly score can be used to detect significant changes. The main result is that they detected anomalies in the graph embeddings three days before the crash; however, the work used a unique graph, without distinguishing the stablecoins, hence the detected anomalies before it may not necessarily be related to the Luna Terra system or the crash. Similarly, Zhu et al. [31], focuses on the task of trend detection in blockchain-based networks, with a method focused on core decomposition to detect a change in the activity in the temporal transaction network. Among the selected datasets for evaluation, they rely on a single dataset: the authors can detect a perturbation in network structure caused by the crash, but they too model the entire transaction data as a homogeneous temporal network.
@@ 3 Research motivation
Are the events preceding it unusual in scale or characteristics? Are they isolated to just one currency? We look at the volume and number of trades in each currency over the immediate time period before the attack to find those unusual “whale" events that precede the crash and the aftershocks from the event.
@@
4 Methodology
In this section we give a basic description of our dataset and the methodology we will use for analysis. It should be remembered throughout this and the next section that the analysis of USTC and WLUNC are looking only at those transactions that can be seen on the Ethereum blockchain. The number of transactions on the Terra blockchain (that we cannot observe) where Terra and Luna are traded natively are likely to be far larger in volume.
4.1 Dataset
The dataset we rely on is the stablecoins dataset presented in Chartalist datasets by Shamsi et al. [4]. It covers the activity of users of the top stablecoins traded on Ethereum ordered by market capitalisation at the time of the crash, alongside USTC and WLUNC. The dataset consists of 70,678,297 transactions, occurring between 1st April 2022 and 1st October 2022, covering a month prior to the crash and six months after the crash. We are focused on specific details, namely the sender ( from_address) and the receiver (to_address) of the currency, the transfer timestamp (time_stamp), the token amount sent (value), and the smart contract address (contract_address) that identifies the currency traded. The currencies under examination are associated with the Terra Ecosystem: USTC and WLUNC [^2] In the dataset, WLUNC is denoted as LUNC, but the contract address points to Wrapped LUNA https://etherscan.io/token/0xd2877702675e6ceb975b4a1dff9fb7baf4c91ea9, currently traded as WLUNC . Additionally, the dataset encompasses user transactions involving other stablecoins, including Tether (USDT), a collateral-backed cryptocurrency stablecoin [32]; Dai, a USD-pegged stablecoin maintained through smart contracts, governed by MakerDAO [33]; Pax Dollar (USDP), a stablecoin collateralized 1:1 by USD in Paxos-owned US bank accounts and approved by Wall Street regulators [34]; and USD Coin (USDC), managed by the Centre consortium, backed by a dollars in reserve or other undisclosed "approved investments" [35]. We provided a summary of the currencies in Table 1.
We also integrate the available data with currency daily prices. We recover the price data from the web platform CoinCodex [36]. The obtained data is presented in Figure 1. From Figure 1 we can see how Luna was extremely valuable with a peak value surpassing \$120 a token. Figure 1 presents an alternative version of the plot that gives us a better view of the stablecoins. The plot is normalized by the starting value observed in the period, highlighting the fall of USTC while other stablecoins largely exhibit their designed behaviour, holding a value very close to one dollar.

### v2

# Investigating the Luna–Terra collapse through the temporal multilayer graph structure of the ethereum stablecoin ecosystem

@@
Abstract.
Blockchain technology and cryptocurrencies have garnered considerable attention over the past fifteen years. The term Web3 (sometimes Web 3.0) has been coined to define a possible direction for the web based on the use of decentralisation via blockchain. Cryptocurrencies are characterised by high market volatility and susceptibility to substantial crashes, issues that require temporal analysis methodologies able to tackle the high temporal resolution, heterogeneity and scale of blockchain data. While existing research attempts to analyse crash events, fundamental questions persist regarding the optimal time scale for analysis, differentiation between long-term and short-term trends, and the identification and characterisation of shock events within these decentralised systems.
@@ Abstract.
This paper is novel in its use of temporal, cross-chain graph analysis to explore a cryptocurrency collapse. It emphasises the importance of temporal analysis for studies on web-derived data. In addition, the methodology shows how graph-based analysis can enhance traditional econometric results. Overall, this research carries implications beyond its field, for example for regulatory agencies aiming to safeguard users could use multi-layer temporal graphs as part of their suite of analysis tools.
transaction network, temporal network, cryptocurrency, Web3, multilayer graph
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXcopyright: noneccs: Mathematics of computing Time series analysisccs: Networks Network economicsccs: Applied computing Economicsccs: Information systems Digital cashccs: Information systems Web miningccs: Networks Network dynamics
1. Introduction
@@
Analysing multi-layer, temporal networks at scale is an important problem in the study of complex networks. In particular, there are issues of understanding the best time-scale to analyse, how to determine long-term and short-term trends and how to recognise and characterise shock events in such a system. In this paper, we look at six cryptocurrencies traded over the Ethereum blockchain. We use multi-layer, temporal network analysis techniques which allow us to simultaneously look at how temporal networks evolve in time and at different time-scales while considering several network layers together.
@@
Digital currencies remain an important but potentially volatile market. In particular, there is the possibility that a particular currency will completely crash, wiping the value of that coin to a tiny fraction of its previous trade value: these events can destroy investments and savings, impacting the lives of investors. It is, therefore, important to analyse the events that lead up to a crash and to look at whether the effects of a crash can destabilise other currencies. “Stablecoins” are an attempt to make cryptocurrencies that can be used for trading rather than speculation by “pegging” the value of the coin to a fixed exchange rate (often one dollar to one stablecoin). The stablecoin TerraUSD (UST) [^1] Throughout this paper we will refer to currency by the shortened name it was commonly traded under. was pegged to trade for \$1 and was associated with a companion currency LUNA that was intended to stabilise it. Over a few days in May 2022 both UST and LUNA had collapsed and never regained their value. The collapse is believed to have been precipitated by deliberate and coordinated large-scale sales of UST in an attempt to crash UST, drive down the price of Bitcoin (BTC) and derive profit from this. Some authors (D. et al., 2023; Conlon et al., 2023) claim that it was the collapse of the Luna–Terra ecosystem that led to the well-known failure of FTX.
@@
In this paper, we use techniques from complex network analysis and, in particular, temporal multi-layer analysis to look at the events leading up to and immediately following the collapse of LUNA and UST. UST and LUNA were traded on a cryptocurrency platform known as Terra and this data is not available to us. However, the two currencies were also traded in a “wrapped” form on the Ethereum platform, USTC (US Terra classic) and WLUNC (wrapped Luna classic) respectively. From the available set of Ethereum trades, we look at USTC and WLUNC trades and compare them with the other most highly traded stablecoins on the platform. This gives us a pool of six currencies trading over the same platform that can be compared as they evolve in the weeks before and after the collapse, hence the analysis considers 26,209,860 transactions occurring in a symmetrical analysis period around the day of the crash.
@@
This paper uses temporal, cross-chain, graph analysis to look at the collapse of a cryptocurrency. The analysis in this paper can only be achieved with multi-layer temporal graph systems as it is necessary to look at correlations between different graph layers and to consider the evolution of the system at different time scales. We show that the six currencies are initially highly correlated in terms of the number of sales but, post-crash, separate into two groups where the “failed” currencies correlate with each other but not the other four. We show that two high-volume selling events on UST had a significant effect on the structure of sales volumes before the crash (one event exactly coinciding with the crash) effectively decoupling its sales volume from all other currencies. We also characterise a significant “rally” event that temporarily caused large changes to the UST and LUNA ecosystems shortly after the collapse. The collapse event has significant effects on graph metrics such as graph density and average degree but also effects on measures like clustering coefficient and weakly connected components.
2. Background
@@
Stablecoins are digital currencies that attempt to peg their value to some other reference asset, for example the US dollar or a fixed amount of gold (Ante et al., 2023). This is designed to reduce the volatility of the currency as the high volatility of some currencies can put users off adopting them for use as currency (as opposed to speculation). Many different stablecoins exist traded over several different blockchain systems. Digital currencies can be traded between different blockchains by a process known as “wrapping”. In this process, a currency token originating on chain A is held in place and a wrapped coin representing the same token is minted on chain B (Caldarelli, 2022). This allows, for example, BTC to be spent on the Ethereum chain as Wrapped Bitcoin (WBTC).
@@
Table 1. Currencies under study, the acronym is the label that the currency uses for trading.
In this paper, we wish to study the ecosystem around the currencies Terra and Luna which ran on their own blockchain known as Terra. TerraUSD(UST) was a stablecoin pegged at one US dollar. Luna (LUNA) was a cryptocurrency designed to ensure the stability of the Terra stablecoin. This was achieved by an algorithmic process known as “Anchor”, that automatically would swap Terra with Luna currency to adjust their prices: in fact, the protocol kept the UST exchange rate stable by varying the pool of Terra available (Briola et al., 2023), generating or destroying tokens when needed. The Terra blockchain data is unavailable to us, however, we can study wrapped versions of UST and LUNA traded on the Ethereum blockchain. We provided a summary of the currencies under study in Table 1.
The interest in studying the Terra/Luna ecosystem comes from the sudden complete crash of UST in May 2022. A timeline extracted from (Briola et al., 2023) and (Shamsi et al., 2022) is as follows:
@@
  • 7th May 2022: Agreement is made to buy a large number of UST for BTC greatly reducing the liquidity in the pool. Simultaneously a large number of UST tokens are put up for sale (which caused the first UST de-pegging, below \$0.99 (Briola et al., 2023)).
@@
  • 25th May 2022: Proposal (Luna, 2022; Terra, 2022; MC, 2022) for the launch of a new blockchain Terra 2.0 is official.
  • 27th May 2022 (T2): Terra 2.0 is live. Original LUNA tokens became Luna Classic (LUNC). The original UST stablecoin became TerraClassicUSD (USTC). Unlike the original Terra network, Terra 2.0 has no algorithmic stablecoin (Labs, 2023). We denote this event as Terra 2.0 (T2).
The final consequence was not only the collapse of the Terra platform, UST and LUNA but also a decrease in the price of BTC. The motivation speculated for the attack is that the attackers were short-selling large amounts of BTC. Short selling is, in effect, a bet that a traded asset will decrease in price. The short seller contracts to borrow and sell the asset now and buy it at a later date at market price. If the price drops the seller, has profited by the amount of the price drop. In some markets, short selling is highly regulated as it has been implicated in market crashes (Jiang et al., 2022). In our case study, there is a belief that high levels of selling of Bitcoin observed in the days prior to the crisis were linked to short selling  (Briola et al., 2023).
@@
The Luna crash serves as an interesting case study due to its relevance within the stablecoins ecosystem, attracting attention from both industry and the research community. There are many works describing the ecosystem and the properties of the Terra system. For instance, Uhlig (Uhlig, 2022) proposes a theoretical framework for the analysis of hourly data on prices and market capitalisation data, delineating the gradual unfolding of the crash. Limited research leverages on-chain data. Cho (Cho, 2023) investigates the underlying causes of the de-pegging event focusing on the token economics of the Terra Blockchain. They utilise on-chain data focusing on transactions converting USTC to LUNC over the four days following the May 9 crash, identifying critical design flaws in the conversion process exploited by attackers. These works focus mainly on Terra. However, various studies explore the impact of the crash on other currencies. De Blasis et al. (De Blasis et al., 2023) examine the impact of the USTC collapse on other stablecoins by using proprietary minute-by-minute price transaction data of various cryptocurrencies, over a symmetrical 40-day period around the USTC crash of May 9th. Their work shows the presence of significant contagion from the USTC collapse, which could be explained by traders’ herding behaviour i.e. users reacting to the crash by moving to other currencies. Briola et al. (Briola et al., 2023) provide a background on the Terra project and describe the main events that led to the crash, before delving into an analysis focused on the hourly price data time series. They study the evolution of dependencies between 61 cryptocurrencies, by leveraging the price time series to define a network structure among currencies, a correlation network: their analysis seems to exclude the presence of herding behaviour. The importance of analyzing cross-chain behaviour is denoted by other works as well. Barthere et al. (Barthere et al., 2022) analyses proprietary data from Terra and Ethereum to identify the main users involved in the crash. They assert that a small subset of users initiated the crash by leveraging system vulnerabilities, benefiting significantly by converting USTC to other coins before the crash. Liu et al. (Liu et al., 2023) employ a similar dataset and approach, leveraging proprietary transaction data from Terra and Ethereum. Their extensive analysis across multiple chains and assets reveals the complex nature of the run on Terra. At the same time, they believe the crash was exacerbated by concerns about system sustainability and not just a concentrated market manipulation. Although they have user transaction data, in their study, the authors do not focus on the transaction networks but focus more on describing the events and identifying the main actors. Meanwhile Vidal-Tomás et al. (D. et al., 2023) look at the wider context including the well-known FTX crash from the perspective of the Luna–Terra crash, citing the latter as a significant contributing factor to the former. Conlon et al. (Conlon et al., 2023) take a similar view.
@@
A key limitation in these works is while they provide important results on the crash, they neglect the analysis of the transactions among users from a network standpoint: indeed, the circulation of tokens within the ecosystem can be effectively represented as networks or graphs (Barabási, 2013) a useful tutorial is found in (Khan and Akcora, 2022). In this representation, nodes (or vertices) symbolise users, while links (or edges) denote the dynamic interactions or relationships among them. This modelling and analysis approach for currency systems through networks has significantly enhanced our understanding of financial transaction networks (Aridhi and Nguifo, 2016; Coscia, 2021). Moreover, temporal networks (Holme and Saramäki, 2019) have proven to be a valuable framework for comprehending web-based currency systems such as Bitcoin (Meiklejohn et al., 2013; Wang et al., 2021; Maesa et al., 2018; Gensollen and Latapy, 2020) and Ethereum (Guo et al., 2019; Sharma et al., 2023; Nadini et al., 2021; Galdeman et al., 2022), as well as in the context of new-generation social media platforms (Guidi et al., 2021; Ba et al., 2022, 2023). However, none of these works analyses the stablecoin ecosystem from a network standpoint. Recent works analyzing the Chartalist dataset by Shamsi et al. (Shamsi et al., 2022) highlighted the importance of network structure in the stablecoin system. For instance, Huang et al. (Huang et al., 2023) focus on change detection on dynamic graphs, proposing a spectral method to create dynamic graph embeddings and an associated anomaly score can be used to detect significant changes. The main result is that they detected anomalies in the graph embeddings three days before the crash; however, the work used a unique graph, without distinguishing the stablecoins, hence the detected anomalies before it may not necessarily be related to the Luna Terra system or the crash. Similarly, Zhu et al. (Zhu et al., 2023), focuses on the task of trend detection in blockchain-based networks, with a method focused on core decomposition to detect a change in the activity in the temporal transaction network. Among the selected datasets for evaluation, they rely on a single dataset: the authors can detect a perturbation in network structure caused by the crash, but they too model the entire transaction data as a homogeneous temporal network.
@@ 3. Research motivation
Are the events preceding it unusual in scale or characteristics? Are they isolated to just one currency? We look at the volume and number of trades in each currency over the immediate time period before the attack to find those unusual “whale” events that precede the crash and the aftershocks from the event.
@@
4. Methodology
In this section we give a description of our dataset and the methodology we will use for analysis. It should be remembered throughout this and the next section that the analysis of USTC and WLUNC are looking only at those transactions that can be seen on the Ethereum blockchain. The number of transactions on the Terra blockchain (that we cannot observe) where Terra and Luna are traded natively are likely to be far larger in volume. Our analysis is conducted using the open-source, temporal graph software Raphtory developed at Queen Mary University of London and Pometry Ltd (Steer et al., 2023). The software allows us to conduct analysis across multiple layers, over multiple time windows simultaneously.
4.1. Dataset
The dataset we rely on is the stablecoins dataset presented in Chartalist datasets by Shamsi et al. (Shamsi et al., 2022). It covers the activity of users of the top stablecoins traded on Ethereum ordered by market capitalisation at the time of the crash, alongside USTC and WLUNC. The dataset consists of 70,678,297 transactions, occurring between 1st April 2022 and 1st October 2022, covering a month prior to the crash and six months after the crash. We are focused on specific details, namely the sender ( from_address) and the receiver (to_address) of the currency, the transfer timestamp (time_stamp), the token amount sent (value), and the smart contract address (contract_address) that identifies the currency traded. The currencies under examination are associated with the Terra Ecosystem: USTC and WLUNC [^2] In the dataset, WLUNC is denoted as LUNC, but the contract address points to Wrapped LUNA https://etherscan.io/token/0xd2877702675e6ceb975b4a1dff9fb7baf4c91ea9, currently traded as WLUNC . Additionally, the dataset encompasses user transactions involving other stablecoins, including Tether (USDT), a collateral-backed cryptocurrency stablecoin (Inc., 2022); Dai, a USD-pegged stablecoin maintained through smart contracts, governed by MakerDAO (Team, 2020); Pax Dollar (USDP), a stablecoin collateralized 1:1 by USD in Paxos-owned US bank accounts and approved by Wall Street regulators (Paxos, 2021); and USD Coin (USDC), managed by the Centre consortium, backed by a dollars in reserve or other undisclosed ”approved investments” (Centre, 2018). We provided a summary of the currencies in Table 1.
We also integrate the available data with currency daily prices. We recover the price data from the web platform CoinCodex (CoinCodex, 2023). The obtained data is presented in Figure 1. From Figure 1 we can see how Luna was extremely valuable with a peak value surpassing \$120 a token. Figure 1 presents an alternative version of the plot that gives us a better view of the stablecoins. The plot is normalized by the starting value observed in the period, highlighting the fall of USTC while other stablecoins largely exhibit their designed behaviour, holding a value very close to one dollar.